# The cosmological population of Gamma-ray Bursts from the disks of Active Galactic Nuclei

### v1

@@ 3.2 Results
Figure 10: Cumulative distribution functions for the afterglow flux densities of SGRBs (top row) and LGRB (bottom row) across multiple energy bands: radio (left), optical (second from left), X-ray (second from right), and gamma-rays (right). Similarly to Fig. 9, undiffused (solid line) and diffused (dashed line) distributions are separately shown. Blue and orange lines represent the curves for the SG and TQM models, respectively. The vertical black lines indicate the sensitivity thresholds for specific observatories and integration times (see text): VLA for radio, HST for optical, Chandra for X-ray, and Fermi GBMfor gamma-ray. The highest fraction of observable events is present in X-ray for both SG and TQM models. While SG model has a similar fraction in optical and gamma rays, TQM model has a substantially higher fraction in optical than in gamma rays. The entire intrinsic and diffused populations in radio are below the VLA limit and thus undetectable.
@@
Fig. 9 shows the cumulative probability distribution of observing the prompt emission from a SGRB (left panel) or a long GRB (right panel) in an AGN disk, with flux above a certain value. The threshold corresponding to the sensitivity of the Fermi GBM detector is indicated with a vertical black line as a reference. As expected, the detection probability is very low in the SG diffused model, hovering well below $\sim 1\%$ for both SGRBs and LGRBs. As for TQM, the detectaion probability is $2\sim 3\%$ for both SGRBs and LGRBs, with the latter having the higher probability. The SG model has a slightly higher fraction of detectability if these transients were to emerge undiffused via a lower density funnel produced by their progenitors. Quantitatively, the SG undiffused model expects a fraction $\sim 41\%$ of SGRBs and of $\sim 53\%$ of LGRBs to be detectable in prompt $\gamma$-rays, while TQM undiffused model expects a fraction of $\sim 38\%$ of SGRBs and of $\sim 49\%$ of LGRBs to be detectable.
@@
For the GRBs which are able to emerge, it is interesting to investigate the distribution of the duration times $T_{90}$. Reminding the discussion earlier, Lazzati et al. (2022) identified a critical density (whose exact value depends on the burst energetic and Lorentz factor), above which the external shock forms before the internal one, resulting in a slightly dimmer GRB but with longer duration $T_{90}$. The statistical outcome of the duration distribution is shown in Fig. 16, for both the short GRBs (left panel) and the long ones (right panel). In both cases, the intrinsic distribution of engine durations is compared to that of $T_{90}$. For both the undiffused and the diffused scenarios, we display both the intrinsic distribution as well as the distributions corresponding to the subset of GRBs whose flux is above the detection limit of Fermi. For LGRBs, the peak of the distribution shifts from several tens of seconds to several hundreds of seconds. These could hence be mis-classified as belonging to the class of very long GRBs (Gendre et al., 2013). Conversely, short GRBs, whose intrinsic distribution peaks at a fraction of a second, are found to peak in $T_{90}$ at several seconds, hence making these short GRBs rather look as long GRBs, as argued for the specific case of GRB191019A (Lazzati et al. 2023, but see Stratta et al. 2024).
@@ 4 Summary and Discussion
In the diffused case, GRBs emerging from outer disk regions, where density and optical depth are relatively lower, are more likely to be detectable, whereas those originating in the inner regions face significant suppression. Specifically, we expect to observe GRBs mostly from low redshifts and AGN disks with large SMBH mass, $\gtrsim 10^{7.5}M_{\odot}$ and in outer regions of the disk, $R\sim[10^{6},10^{7}]R_{g}$. Similarly, in the undiffused case, we expect to observe GRBs mostly from low redshifts and AGN disks with large SMBH masses. However, there is also a substantial contribution from intermediate regions of the disk. In particular, the prompt emission has a similar probability to emerge from various regions of the disk regardless of the disk model.
For both disk models, GRBs would most likely be detectable via the prompt $\gamma$-ray emission and the afterglow emission in the X-ray band. The SG model, characterized by regions of higher densities and optical depths, results in suppressed detectability, particularly in the diffused scenario. However, in the undiffused case, the prompt emission dominates as the primary observational window due to the stronger external shocks that enhance synchrotron radiation. On the other hand, the TQM model, with its orders of magnitude lower densities than the SG model, and less extreme optical depths, favors detectability across a wider range of wavelengths, especially in the X-ray band and with some contribution in the optical waveband. These results underscore the crucial dependence of GRB detectability on disk models, where SG disk environments amplify prompt emission through higher densities, whereas TQM environments provide a broader wavelength range of detectable signals due to their relatively lower optical depths. The differences between the two models are so consequential that it might be possible to use even a few GRB detections to learn about the AGN disk structure.
@@
In addition to GRBs, AGN disks are expected to host a variety of electromagnetic transients that could produce overlapping observational signatures. These include tidal disruption events from the central SMBH (Metzger et al., 2012), micro tidal disruption events from stellar-mass BHs (Yang et al., 2022), jets from merging binary BHs (Tagawa et al., 2023a) as well as isolated BHs undergoing hyperaccretion (Pan & Yang, 2021; Tagawa et al., 2023b; Liu et al., 2024), supernovae (Grishin et al., 2021), interacting kilonovae (Ren et al., 2022), and accretion-induced collapse of neutron stars (Perna et al., 2021b; Zhang et al., 2024) and white dwarves (Zhu et al., 2021a). Each of these events could contribute to the overall transient activity observed in AGN environments, and distinguishing between them will require careful analysis of their light curves, spectral signatures, and temporal evolution. As more data becomes available, we expect to refine the criteria for identifying these AGN-hosted GRBs and disentangling them from the broader transient population.

### v2

@@ 3.2 Results
Figure 10: Cumulative distribution functions for the peak afterglow flux densities of SGRBs (top row) and LGRB (bottom row) across multiple energy bands: radio (left), optical (second from left), X-ray (second from right), and gamma-rays (right). Similarly to Fig. 9, undiffused (solid line) and diffused (dashed line) distributions are separately shown. Blue and orange lines represent the curves for the SG and TQM models, respectively. The vertical black lines indicate the sensitivity thresholds for specific observatories and integration times (see text): VLA for radio, HST for optical, Chandra for X-ray, and Fermi GBM for gamma-ray. The highest fraction of observable events is present in X-ray for both SG and TQM models. While SG model has a similar fraction in optical and gamma rays, TQM model has a substantially higher fraction in optical than in gamma rays. The entire intrinsic and diffused populations in radio are below the VLA limit and thus undetectable.
@@
Fig. 9 shows the cumulative probability distribution of observing the prompt emission from a SGRB (left panel) or a long GRB (right panel) in an AGN disk, with flux above a certain value. The threshold corresponding to the sensitivity of the Fermi GBM detector is indicated with a vertical black line as a reference. As expected, the detection probability is very low in the SG diffused model, hovering well below $\sim 1\%$ for both SGRBs and LGRBs. As for TQM, the detection probability is $2\sim 3\%$ for both SGRBs and LGRBs, with the latter having the higher probability. The SG model has a slightly higher fraction of detectability if these transients were to emerge undiffused via a lower density funnel produced by their progenitors. Quantitatively, the SG undiffused model expects a fraction $\sim 41\%$ of SGRBs and of $\sim 53\%$ of LGRBs to be detectable in prompt $\gamma$-rays, while TQM undiffused model expects a fraction of $\sim 38\%$ of SGRBs and of $\sim 49\%$ of LGRBs to be detectable.
@@
For the GRBs which are able to emerge, it is interesting to investigate the distribution of the duration times $T_{90}$. Reminding the discussion earlier, Lazzati et al. (2022) identified a critical density (whose exact value depends on the burst energetic and Lorentz factor), above which the external shock forms before the internal one, resulting in a slightly dimmer GRB but with longer duration $T_{90}$. The statistical outcome of the duration distribution is shown in Fig. 16, for both the short GRBs (left panel) and the long ones (right panel). In both cases, the intrinsic distribution of engine durations is compared to that of $T_{90}$. For both the undiffused and the diffused scenarios, we display both the intrinsic distribution as well as the distributions corresponding to the subset of GRBs whose flux is above the detection limit of Fermi. For LGRBs, the peak of the distribution shifts from several tens of seconds to several hundreds of seconds. These could hence be mis-classified as belonging to the class of very long GRBs (Gendre et al., 2013). Conversely, short GRBs, whose intrinsic distribution peaks at a fraction of a second, are found to peak in $T_{90}$ at several seconds, hence making these short GRBs rather look as long GRBs, as argued for the specific case of GRB191019A (Lazzati et al. 2023, but see Stratta et al. 2024). Further confirmation for an AGN-disk origin can be provided by longer wavelength observations aimed at detecting the afterglow: counterparts in X-rays are expected if there was a $\gamma$-ray detection. Optical counterparts may be observed with $\sim 40-60\%$ probability only in the most massive disks (due to the lower density), while a radio counterpart is extremely unlikely.
@@ 4 Summary and Discussion
In the diffused case, GRBs emerging from outer disk regions, where density and optical depth are relatively lower, are more likely to be detectable, whereas those originating in the inner regions face significant suppression. Specifically, we expect to observe GRBs mostly from low redshifts and AGN disks with large SMBH mass, $\gtrsim 10^{7.5}M_{\odot}$ and in outer regions of the disk, $R\sim[10^{6},10^{7}]R_{g}$. Similarly, in the undiffused case, we expect to observe GRBs mostly from low redshifts and AGN disks with large SMBH masses. However, there is also a substantial contribution from intermediate regions of the disk.
For both disk models, GRBs would most likely be detectable via the prompt $\gamma$-ray emission and the afterglow emission in the X-ray band. The SG model, characterized by regions of higher densities and optical depths, results in suppressed detectability, particularly in the diffused scenario. However, in the undiffused case, the prompt emission dominates as the primary observational window due to the stronger external shocks that enhance synchrotron radiation. On the other hand, the TQM model, with its orders of magnitude lower densities than the SG model, and less extreme optical depths, favors detectability across a wider range of wavelengths, especially in the X-ray band and with some contribution in the optical waveband. These results underscore the crucial dependence of GRB detectability on disk models, where SG disk environments amplify prompt emission through higher densities, whereas TQM environments provide a broader wavelength range of detectable signals due to their relatively lower optical depths. The differences between the two models are so consequential that it might be possible to use even a few GRB detections to learn about the AGN disk structure. Our results are in broad agreement with those of Zhang et al. (2024a) in that observability is favored for more energetic GRBs, and that there is a strong dependence on the disk properties. However, the specific details vary since they explored the hydrodynamic properties of the jets in dense environments, while we focused on the emission properties and line-of-sight scattering opacity.
@@
In addition to GRBs, AGN disks are expected to host a variety of electromagnetic transients that could produce overlapping observational signatures. These include tidal disruption events from the central SMBH (Metzger et al., 2012), micro tidal disruption events from stellar-mass BHs (Yang et al., 2022), jets from merging binary BHs (Tagawa et al., 2023a) as well as isolated BHs undergoing hyperaccretion (Pan & Yang, 2021; Tagawa et al., 2023b; Liu et al., 2024), supernovae (Grishin et al., 2021), interacting kilonovae (Ren et al., 2022), and accretion-induced collapse of neutron stars (Perna et al., 2021b; Zhang et al., 2024b) and white dwarves (Zhu et al., 2021a). Each of these events could contribute to the overall transient activity observed in AGN environments, and distinguishing between them will require careful analysis of their light curves, spectral signatures, and temporal evolution. As more data becomes available, we expect to refine the criteria for identifying these AGN-hosted GRBs and disentangling them from the broader transient population.